OpenStudy (anonymous):

solve. 60 + 30w=480 A. 16 B.13 C. 15 D. 14

OpenStudy (hoblos):

subtract 60 then divide by 30

OpenStudy (hoblos):

60 + 30w=480 30w = 420 w=14

OpenStudy (anonymous):

60 + 30w = 480 -60 -60 30w = 420 /30 /30 w = 14